OpenStudy (anonymous):

Which table best represents the steps to solve the equation 8 + x over 3 = 4?

OpenStudy (anonymous):

tep Justification Step 1: x over 3 = 4 + 8 The same number +8 can be added to both sides of the equation Step 2: x over 3 = 12 Adding 8 and 4 on the right hand side Step 3: x = 12 • 3 The same number 3 can be multiplied on both sides of the equation Step 4: x = 36 Multiplying 12 and 3 on the right hand side Step Justification Step 1: x over 3 = 4 - 8 The same number -8 can be added to both sides of the equation Step 2: x over 3 = -4 Adding -8 and 4 on the right hand side Step 3: x = -4 • 3 The same number 3 can be multiplied on both sides of the equation Step 4: x = -12 Multiplying -4 and 3 on the right hand side Step Justification Step 1: x over 3 = 4 - 8 The same number - 8 can be added to both sides of the equation Step 2: x over 3 = -4 Adding 4 and -8 on the right hand side tep 3: x = - 4 - 3 The same number 3 can be subtracted from both sides of the equation Step 4: x = -7 Simplifying -4 -3 Step Justification Step 1: x over 3 = -4 - 8 The same number -8 can be added to both sides of the equation Step 2: x over 3 = -4 Adding -8 and 4 on the right hand side Step 3: x = -4 ÷ 3 The same number 3 can be divided on both sides of the equation Step 4: x = -1.33 Dividing -4 by 3 on the right hand side

OpenStudy (imstuck):

Let me work it for you step by step and you are sure to not only figure out the right answer, but also be able to do them on your own!